OpenStudy (anonymous):

Solve the system of linear equations below. Enter your solution as an ordered pair (x,y). Do not include spaces in your answer. 2x+2y =2 x - y =5

OpenStudy (rock_mit182):

\[(1) 2x+2y =2\] \[(2) x-y =5\]

OpenStudy (rock_mit182):

now we're going to solve this by reduction method

OpenStudy (rock_mit182):

multiply the equation (2) by 2 and then add (1) and (2)

OpenStudy (rock_mit182):

to get rid off the y

OpenStudy (rock_mit182):

@Gusanocx

OpenStudy (anonymous):

Its kinda confusing I just learned this but I kinda forget the steps in this.

OpenStudy (rock_mit182):

show me some work i will help you out

OpenStudy (rock_mit182):

first of all, follow the step that i told you the other part is so easy

OpenStudy (anonymous):

So I multiply the equation (2) by that is x-y=5 so is it x*y then you said I add both equations.

OpenStudy (anonymous):

^2

OpenStudy (rock_mit182):

2 times the eqution (2)

OpenStudy (rock_mit182):

so we left with :\[2x-2y=10\]

OpenStudy (rock_mit182):

that makes sense for you ?

OpenStudy (anonymous):

2(2+2y) + y then you said I'm left with 2x-2y=10

OpenStudy (anonymous):

or in the parentheses is it a -

OpenStudy (anonymous):

(3,-2)

OpenStudy (rock_mit182):

yeah now add the equation (1) with (2)

OpenStudy (rock_mit182):

then solve for x

OpenStudy (rock_mit182):

and finally plug and chuff that value, in the equation (1) or (2) to find the value of Y

OpenStudy (anonymous):

thank you (:

OpenStudy (rock_mit182):

you're welcome

OpenStudy (rock_mit182):

;)